OpenStudy (anonymous):

What is an asymptote of the hyperbola given by the equation y^2/1 - x^2/81 = 1?

OpenStudy (anonymous):

Find the square roots of 1 and 81. THen put them in y/x \[y = ±\sqrt{\frac{1}{81}}\]